# dMel: Speech Tokenization made Simple

### v1

@@ 3.2 Training Configuration
We train decoder-only transformers in three different sizes: Small, Base, and Large (see Appendix Table 9). By default, the Base model is used in all experiments if not stated otherwise. All models use pre-LayerNorm with dropout set to 0.1 for residual, attention and embedding and 0.3 for positional embedding. dMel uses 16 discrete bins for each channel while text is tokenized with a character vocabulary; the speaker embedding dvector has 512 dimensions(see Appendx D for details).
@@ 3.3 Main Results
In this section, we first evaluate different tokenization methods on speech reconstruction. Then we report the TTS and ASR results using an LM-style (decoder-only) model with different speech tokens.
@@ 3.3.1 Speech Reconstruction
Following [40], we randomly sample 300 speech utterances and their ground truth transcriptions from the LibriSpeech test-clean dataset. We use the speech2unit and unit2speech modules to convert the speech signal to speech tokens and then reconstruct the speech signal from the speech tokens. We compute the WER between the ASR outputs from HuBERT-Large [14] [^4] We use checkpoint https://huggingface.co/facebook/hubert-large-ls960-ft%****␣experiments.tex␣Line␣25␣****. on the audio samples and their ground truth transcripts. We also report MOS-LQO (Mean Opinion Score - Listening Quality Objective) score to measure the reconstruction quality using ViSQOL [13]. Finally, use human evaluation to measure the naturalness of the reconstructed speech using a MOS score. We instruct the human evaluators to rate the naturalness of the reconstructed speech on a scale of 1 to 5, where 1 is the worst and 5 is the best. The results are shown in Table 3.
Table 3: Speech reconstruction results on 300 random samples from LibriSpeech test-clean set. For each tokenizer, we list its frame rate and the model parameter sizes of its speech2unit and unit2speech modules. WER is evaluated with HuBERT-Large and MOS-LQO with ViSQOL.
Tokenizer Speech2unit Unit2speech Frame Rate WER MOS-LQO MOS
GroundTruth - - - 2.02 - 3.91
HuBERT-KM [19] 95M  [^1] M 50Hz 8.71 2.06 2.78
EnCodec [11] 7M 7M 75Hz 2.03 4.03 3.77
SpeechTokenizer [40] 65M 34M 50Hz 2.41 4.19 3.89
Mel-HifiGAN [18] n/a 12M 80Hz 2.08 4.52 3.85
dMel-HifiGAN n/a 12M 80Hz 2.11 4.47 3.73
Mel-PWG [36] n/a 1M 80Hz 2.13 4.40 3.19
dMel-PWG n/a 1M 80Hz 2.23 4.37 3.07
Mel-PWG n/a 1M 40Hz 2.36 4.34 2.90
dMel-PWG n/a 1M 40Hz 2.51 4.29 2.76
@@
We apply different vocoders to reconstruct the speech signal from mel-filterbanks, and find that the WER of the reconstructed speech signal is comparable to the acoustic tokenization methods with a fraction of the parameters. Also, mel-filterbanks achieve a better MOS-LQO score, which indicates that the reconstructed audio is more similar to the original audio. By comparing Mel and dMel, we can see that discretization has little impact on WER and MOS-LQO scores. We also find that the exact vocoder matters much less than the frame rate of tokenization: the WER goes from 2.08 to 2.13 when switching from HifiGAN to ParallelWaveGAN, but it falls from 2.13 to 2.36 when the frame rate is changed from 80Hz to 40Hz. However, even a 1M parameter vocoder operating at a 40Hz frame rate is comparable to the much larger SpeechTokenizer on WER and MOS-LQO metrics.
@@ 3.3.2 LM-Style Text-to-Speech
As shown in Table 4 for training on LibriSpeech dataset, our decoder-only model with dMel tokenization achieves a WER of 4.3 and a CER of 1.8, significantly outperforming the baseline methods. This indicates that our model can generate more accurate speech with less hallucination and distortion. Furthermore, we observed that the AR model trained on SpeechTokenizer-trained model exhibits a much higher WER compared to the idiosyncratic coarse to fine models (labeled AR+NAR) developed for these residual tokenizers – indicating that dMel lies on a simpler data manifold.
Given the success of our decoder-only dMel TTS model, dubbed RichTTS, we further evaluate it on various datasets, including LJSpeech [10], VCTK [35], and LibriTTS [38], and compare it with popular TTS models, including Tacotron2 [28], FastSpeech2 [26], and VITS [16]. We conduct human evaluation to measure the naturalness of 50 randomly sampled synthesized speech from VCTK test set. The results are shown in Table 5. Our model achieves competitive performance on the TTS task in terms of both MOS and WER, demonstrating its effectiveness in generating high-quality synthesized speech. Interestingly, we find that VITS performs poorly on the VCTK WER. We suspect that this is because VITS tends to make more mistakes at the beginning of each sequence, and since VCTK comprises short sequences, even one or two word errors can lead to a high WER.
Table 4: Text-to-speech results for different tokenizers. All models are trained on LibriSpeech 960h dataset. Evaluation is done via speech generation on the full test-clean transcriptions and speakers, and then evaluated WER with WhisperX base.en. We use †  to annotate results from the reference paper. RichTTS is our model trained with dMel tokenization.
WER (%) CER (%)
RichTTS (HuBERT+KM), 258M 9.5 4.3
VOXTLM (HuBERT+KM)† [21], 350M - 3.5
RichTTS (SpeechTokenizer), AR, 258M 11.4 5.9
USLM (SpeechTokenizer)†, AR+NAR [40], 356M 6.5 -
RichTTS (dMel), 258M 4.3 1.8
Table 5: WER (%) (evaluated with WhisperX base.en ASR model) of different TTS models’ generations using transcriptions from the evaluation set of each dataset. Each column corresponds to the training and evaluation dataset. We used ESP-Net [33] to generate the results for prior work.
WER MOS
LJSpeech VCTK LibriTTS VCTK
Tacotron2 [28] 4.4 4.9 7.3 1.3
FastSpeech2 [26] 6.1 8.7 10.2 1.3
VITS [9] 6.4 24.1 8.3 4.1
RichTTS (Ours) 4.0 2.2 4.5 3.9
@@
Table 6: Text-to-speech results for models trained on Librispeech 960h data and evaluated on the transcriptions from LJSpeech test set. We show WER (%) evaluated by WhisperX base.en ASR model on text grouped by sequence length.
WER Tacotron2 [28] FastSpeech2 [26] VITS(YourTTS) [9] RichTTS (Ours)
total 4.4 6.1 6.4 4.0
10-20 words 5.5 3.1 7.4 3.5
more than 20 words 3.3 9.1 5.3 3.0
@@ 3.3.3 LM-Style Speech-to-Text
Table 7: Our Base ASR (258M) models are trained on LibriSpeech 960h and evaluated for different speech tokenizers (greedy decoding is reported). We compare also with prior SOTA work on decoder-only model for ASR (VOXTLM). Results are shown across 2 runs with mean WER and standard deviation.
dev-clean dev-other test-clean test-other
VOXTLM [21] (HuBERT+KM [19]), 350M - - 6.5 17.6
VOXTLM [21] (HuBERT+KM [19]), 1.3B - - 4.6 12.1
RichASR (SpeechTokenizer [40]) 6.5$\pm$0.3 16.9$\pm$0.7 6.9$\pm$0.4 17.5$\pm$0.5
RichASR (HuBERT+KM [19]) 5.3$\pm$0.1 13.7$\pm$0.2 5.8$\pm$0.1 13.8$\pm$0.1
RichASR (dMel) 3.8$\pm$0.1 10.3$\pm$0.1 4.2$\pm$0.2 10.4$\pm$0.1
@@
Training an LM-style speech-to-text (ASR) model can test if the speech tokens can preserve the semantic information in the speech signal and support the speech understanding task. Table 7 shows results of our model RichASR trained with different tokenizations including dMel for the ASR task. Our decoder-only model with dMel speech tokenization achieves 4.2% WER on the test-clean and 10.4% WER on the test-other sets outperforming both HuBERT-KM and SpeechTokenizer. We also observe that our model with HuBERT-KM [19] outperforms the SpeechTokenizer [40] for ASR, which is reasonable as semantic tokens are more suitable for the ASR task. Additionally, our method outperforms VOXTLM [21], which is a larger model trained with more data and initialized from a pretrained LLM (OPT) using HuBERT-KM (we use exactly the same tokenization with 200 clusters) as the speech tokenizer. The ASR results clearly demonstrate the benefit of using our speech tokenizer for the speech understanding task, as it better preserves the semantic information in the speech signal. Further details and ablations on data size and model size can be found in Appendix D and Appendix E Table 11.
3.3.4 Joint Speech-Text Modeling
Table 8: Results of ASR and TTS jointly trained model on LibriSpeech dataset.
ASR TTS
test-clean test-other WER CER
@@ 3.3.4 Joint Speech-Text Modeling
Table 8 shows that the joint model is worse on both tasks, but ASR is affected more than TTS. Comparing our results to VOXTLM [21], which initializes its model from pretrained LLM (OPT) and finetunes it with multiple tasks and datasets, we speculate that our joint model needs text-only training to learn a good LM for better ASR performance. Our model structure trivially allows for this text-only training, but we leave those experiments for future work.
4 Related Work
Speech Tokenization
Speech tokenization is a critical step in speech processing tasks, enabling the discretization of continuous speech signals into tokens for language model-based processing. Existing speech tokenization methods can be broadly categorized into two types: semantic tokens and acoustic tokens. Semantic tokens are extracted from self-supervised pretrained speech models, where the speech signal is first encoded into speech representations and then clustered into semantic tokens with $k$-means clustering [2, 14]. Acoustic tokens are obtained from pretrained audio compression models, which compress the speech signal into codebook indices with residual vector quantization (RVQ) and reconstruction objectives [37, 11]. To combine the advantage of both semantic and acoustic tokens, AudioLM [7] proposed to modeling both semantic tokens and acoustic tokens with 3 stages: semantic modeling, coarse acoustic modeling, and fine acoustic modeling. This solution covers both the content and speech quality, but its multi-stage hierarchical structure complicates the model and slows the training and inference. AudioPalm [27] follows a similar approach, and show that the scale of speech tokenizer’s training data and model parameters are critical for this multi-stage modeling. This observation indicates the speech compression model is far from lossless, and not generalizable to low-resource scenarios. Another solution is combining the semantic and acoustic feature together, and formulate a new token that capture both the semantic and acoustic information. Zhang et al. [40] proposed to distill the semantic tokens into the acoustic token’s first residual channel during the training of the RVQ model in a teacher-student manner. However, such combined methods are complex and require additional pretraining, complicating the model architecture and increasing computational overhead. Also, the model architecture is still not a single stage model, where only the semantic modeling is skipped. In comparison, dMel is a train-free speech tokenization method that discretizes mel-filterbanks directly into bins, inherently preserving both semantic and acoustic information in a unified representation.
@@ 4 Related Work
Modeling speech and text jointly is a challenging task, as speech signals are continuous and while text is discrete. Existing works have explored various approaches to address this challenge, including usage of separate encoders for different modalities [1, 5]. Bai et al. [3] proposed an encoder only model A3T for speech-text modeling, by introducing alignment embedding to encourage cross-modal transfer between text and speech. Although A3T achieved good performance on speech synthesis and editing tasks, it cannot generate text and cannot generalize to longform generation because of its encoder-only architecture and mask-reconstruction training strategy. VioLA [32] also targets a unified speech-text model which can generate speech and text with a single model, but it is specifically designed for the Encodec [11] style feature, and compelled to model speech tokens in a multi-stage hierarchical manner. Maiti et al. [21] proposed a decoder-only language model VOXTLM, to model speech and text jointly. However, VOXTLM is only models the HuBERT semantic tokens, and relies on an external generation model to transform semantic tokens into waveform, but the speaker and acoustic information are lost. In comparison, the model architecture in this paper is a simple, single stage decoder-only transformer language model, and can handle both the speech generation and text generation tasks.

### v2

@@ 3.2 Training Configuration
We train the LM-style transformers in three different sizes: Small, Base, and Large (see Appendix Table 12). Unless stated otherwise, the Base model is used in all experiments if not stated otherwise. All models use pre-LayerNorm with dropout set to 0.1 for residual, attention and embedding layers and 0.3 for positional embedding. dMel uses 16 discrete bins for each channel while text is tokenized with a character vocabulary; the speaker embedding dvector has 512 dimensions (see Appendx E for details). In all experiments, training data are sampled to 16kHz.
@@ 3.3.1 Speech Reconstruction
Following [46], we randomly sample 300 speech utterances and their ground truth transcriptions from the LibriSpeech test-clean dataset. We use the speech2unit and unit2speech modules to convert the speech signal to speech tokens and then reconstruct the speech signal from the speech tokens. We compute the WER between the ASR outputs from HuBERT-Large [16] [^5] We use checkpoint https://huggingface.co/facebook/hubert-large-ls960-ft. on the audio samples and their ground truth transcripts. We also report MOS-LQO (Mean Opinion Score – Listening Quality Objective) score to measure the reconstruction quality using ViSQOL [15]. Finally, we use human evaluation to measure the naturalness of the reconstructed speech using a MOS score with 95% confidence interval. We instruct the human evaluators to rate the naturalness of the reconstructed speech on a scale of 1 to 5, where 1 is the worst and 5 is the best. The results are shown in Table 3.
Table 3: Speech reconstruction results on 300 random samples from LibriSpeech test-clean set.
Tokenizer
Speech2Unit
(M params)
Unit2Speech
(M params)
Frame
Rate
WER↓ MOS-LQO↑
MOS↑
(95% CI)
GroundTruth - - - 2.02 - 3.91$\pm$0.12
HuBERT-KM 95 111 50Hz 8.71 2.06 2.74$\pm$0.14
EnCodec 7 7 75Hz 2.03 4.03 3.69$\pm$0.13
SpeechTokenizer 65 34 50Hz 2.41 4.19 3.77$\pm$0.13
Mel-HifiGAN n/a 12 80Hz 2.08 4.52 3.80$\pm$0.12
dMel-HifiGAN n/a 12 80Hz 2.11 4.47 3.68$\pm$0.13
Mel-PWG n/a 1 80Hz 2.13 4.40 3.27$\pm$0.14
dMel-PWG n/a 1 80Hz 2.23 4.37 3.23$\pm$0.14
Mel-PWG n/a 1 40Hz 2.36 4.34 2.99$\pm$0.15
dMel-PWG n/a 1 40Hz 2.51 4.29 2.97$\pm$0.15
@@
We apply different vocoders to reconstruct the speech signal from log mel-filterbanks, and find that the WER of the reconstructed speech signal is comparable to the acoustic tokenization methods with a fraction of the parameters. Also, log mel-filterbanks achieve a better MOS-LQO score, which indicates that the reconstructed audio is more similar to the original audio. By comparing Mel and dMel, we can see that discretization has little impact on WER and MOS-LQO scores. We also find that the exact vocoder matters much less than the frame rate of tokenization: the WER goes from 2.08 to 2.13 when switching from HifiGAN to ParallelWaveGAN, but it falls from 2.13 to 2.36 when the frame rate is changed from 80Hz to 40Hz. However, even a 1M parameter vocoder operating at a 40Hz frame rate is comparable to the much larger SpeechTokenizer on WER and MOS-LQO metrics.
@@ 3.3.2 LM-Style Text-to-Speech
As shown in Table 4 for training on LibriSpeech dataset, our LM-style model with dMel tokenization achieves a WER of 4.3 and a CER of 1.8, significantly outperforming the baseline methods. This indicates that our model can generate more accurate speech with less hallucination and distortion. Furthermore, we observed that the AR model trained on SpeechTokenizer tokens exhibits a much higher WER compared to the idiosyncratic coarse to fine models (labeled AR+NAR) developed for these residual tokenizers – indicating that dMel lies on a simpler data manifold.
Given the success of our LM-style dMel TTS model, dubbed RichTTS, we further evaluate it on various datasets, including LJSpeech, VCTK, and LibriTTS, and compare it with popular open-sourced TTS models, including Tacotron2 [33], FastSpeech2 [30], and VITS [18]. We conduct human evaluation to measure the naturalness of 50 randomly sampled synthesized speech from VCTK test set. RichTTS achieves competitive performance on the TTS task in terms of both MOS and WER demonstrating its effectiveness in generating high-quality synthesized speech, see Table 5. Interestingly, we find that VITS performs poorly on the VCTK WER. We suspect this is because VITS tends to make more mistakes at the beginning of each sequence, and since VCTK comprises short sequences, even one or two word errors can lead to a high WER.
Table 4: Text-to-speech results for different tokenizers. RichTTS is trained on LibriSpeech 960h. WER (%) and CER (%) are evaluated with WhisperX ASR (“base.en”) and reported on test-clean.
Model WER↓ (%) CER↓ (%) Params
VOXTLM (HuBERT+KM)†, [24] - 3.5 350M
USLM (SpeechTokenizer)†, AR+NAR, [46] 6.5 - 356M
RichTTS (HuBERT+KM) 9.5 4.3 258M
RichTTS (SpeechTokenizer), AR 11.4 5.9 258M
RichTTS (dMel) 4.3 1.8 258M
Table 5: WER (%) (evaluated with WhisperX ASR “base.en”) and MOS of different TTS models’ generations using transcriptions from each evaluation set that correponds to data used for training.
WER↓ (%)
MOS↑ (95% CI)
Model LJSpeech LibriTTS VCTK VCTK
GroundTruth 2.6 3.8 3.4 4.18$\pm$0.10
Tacotron2, [33] 4.4 7.3 4.2 2.91$\pm$0.15
FastSpeech2, [30] 6.1 10.2 3.8 3.03$\pm$0.14
VITS, [9] 6.4 8.3 11.1 3.56$\pm$0.12
RichTTS (dMel) 4.0 4.5 2.2 3.34$\pm$0.14
Table 6: results for TTS models trained on LibriSpeech 960h and evaluated on LJSpeech test set.
Sequence Length Tacotron2 FastSpeech2 VITS RichTTS
Total WER↓ (%) 4.4 6.1 6.4 4.0
10-20 words 5.5 3.1 7.4 3.5
20+ words 3.3 9.1 5.3 3.0
@@ 3.3.3 LM-Style Speech-to-Text
Table 7: Speech recognition results for different tokenizers measured with WER (%). All models are trained on LibriSpeech 960h.
Model dev-clean↓ dev-other↓ test-clean↓ test-other↓ Params
RichASR (SpeechTokenizer) 6.5$\pm$0.3 16.9$\pm$0.7 6.9$\pm$0.4 17.5$\pm$0.5 258M
RichASR (HuBERT+KM) 5.3$\pm$0.1 13.7$\pm$0.2 5.8$\pm$0.1 13.8$\pm$0.1 258M
RichASR (dMel) 3.8$\pm$0.1 10.3$\pm$0.1 4.2$\pm$0.2 10.4$\pm$0.1 258M
Table 8: Comparison of WER (%) for best RichASR trained with dMel tokenization and prior work with LM-style ASR models and HuBERT+KM with subword modeling on top as tokenization.
Model Data (h) dev-clean↓ dev-other↓ test-clean↓ test-other↓ Params
VOXTLM 280k - - 6.5 17.6 350M
VOXTLM 280k - - 4.6 12.1 1.3B
[10] 960 3.6 7.8 3.8 8.3 355M
RichASR (dMel) 960 3.1 8.4 3.4 8.6 355M
@@
Training an LM-style speech-to-text (ASR) model can test if the speech tokens can preserve the semantic information in the speech signal and support the speech content-based task. Table 7 shows results of our model dubbed RichASR, trained with different tokenizations including dMel for the ASR task. Our LM-style model with dMel speech tokenization achieves 4.2% WER on the test-clean and 10.4% WER on the test-other sets outperforming both HuBERT-KM and SpeechTokenizer. We also observe that our model with HuBERT-KM [22] outperforms the SpeechTokenizer [46] for ASR, which is reasonable as semantic tokens are more suitable for the ASR task.
In Table 8, we further compare RichASR with dMel speech tokenizer trained with GPT-2-meduim architecture [27] on LibriSpeech 960h with prior work: VOXTLM [24] that uses larger model trained with more data and initialized from a pretrained LLM (OPT [45]), and HuBERT-KM with additional subword modeling on top as the speech tokenizer; [10] that also uses GPT-2 architecture trained on LibriSpeech 960h and HuBERT-KM with additional subword modeling on top as the speech tokenizer [^6] We use official codebase to train this model w/o text pretraining as [10] report results only with text pretraining.. RichASR with dMel outperforms VOXTLM; it also outperforms [10] on clean sets and a bit behind it on other sets.
The ASR results clearly demonstrate the benefit of using our dMel speech tokenizer for the content-related tasks in speech, as it better preserves the semantic information in the speech signal. Further details and ablations can be found in Appendix E and F.
3.3.4 Ablations
We first investigate the impact of the codebook sizes, shown in Table 9. The 16-bin configuration used in the paper demonstrates the best overall performance across tasks. While the 32-bin setup slightly outperforms on the ASR test-other set, it shows degraded performance in TTS. This trade-off likely stems from the increased speech vocabulary size, which may pose challenges for accurate prediction. The results may get better with increased data and model size. And 8-bin configuration looses too much information with discretization.
We then ablate the ASR results to understand why ASR LM-style model is behind the state-of-the-art ASR results on LibriSpeech. We take two existing transformer ASR baselines, Seq2Seq and CTC, that use 80 log mel-filterbanks and characters as targets. We then modify these baselines by using dMel instead (the discretization, embedding layer and linear layer) while keeping all other hyper-parameters the same (we adjust only the SpecAugment time masking max width accordingly to keep total masking in ms the same). Our results (Table 10) suggest: i) dMel brings only small degradation compared to Mel; ii) additional discrepancy is coming from different hop distance in featurization; iii) the main and significant performance degradation is coming from switching to LM-style model. The latter is in line with [24] and [10], though was not discussed in detail by any prior work. We hypothesise this gap is due to observed overfitting of the LM-style models.
Table 9: ASR and TTS results (WER, %) with dMel speech tokenizer and different number of bins (codebook size) for discretization in dMel. All models are trained on LibriSpeech 960h.
N-Bins ASR test-clean↓ ASR test-other↓ TTS WER↓ (%)
8 6.6 16.5 7.3
16 4.4 10.7 4.8
32 4.7 10.2 5.7
Table 10: WER (%) comparison for CTC, Seq2Seq, and LM-style ASR models ($\sim$260M) trained on LibriSpeech 960h with dMel and Mel features. We compute 80 log mel-filterbanks with 25ms (50ms) window and 10ms (25ms) hop distance, denoted as ‘10ms’ (‘25ms’).
Model Features dev-clean↓ dev-other↓
[14] (RNN-T – Conformer) Mel-10ms 1.9 4.1
[20] (CTC – Squeezeformer) Mel-10ms 2.3 5.8
Seq2Seq [12] Mel-10ms 2.4 5.4
dMel-10ms 2.5 5.9
Mel-25ms 2.8 6.5
dMel-25ms 2.7 6.2
CTC [13] Mel-10ms 2.1 5.4
dMel-10ms 2.1 5.6
Mel-25ms 2.1 5.4
dMel-25ms 2.3 6.1
LM-style dMel-25ms 3.4 9.5
3.3.5 Unlocking Joint Speech-Text Modeling
Table 11 shows that the joint model is worse on both tasks, but ASR is affected more than TTS. Comparing our results to VOXTLM, which initializes its model from pretrained LLM (OPT) and finetunes it with multiple tasks and datasets, we speculate that our joint model needs text-only training to learn a good LM for better ASR performance. Our model structure trivially allows for this text-only training, but we leave those experiments for future work (for further discussion see Appendix F.3).
4 Related Work
Speech Tokenization
Recent advancements in speech tokenization have primarily focused on two approaches: semantic tokens and acoustic tokens. This section examines these methods, their combinations, and their limitations, highlighting the need for more efficient and generalizable solutions. Semantic tokens, extracted from self-supervised pretrained speech models, have shown promise in capturing high-level content information. Methods like wav2vec [2] and HuBERT [16] employ $k$-means clustering on speech representations to generate these tokens. While effective in capturing semantic content, these approaches often struggle with preserving fine-grained acoustic details crucial for high-quality speech synthesis. In contrast, acoustic tokens, derived from pretrained audio compression models, excel at preserving low-level acoustic information. Techniques such as SoundStream [42] and EnCodec [11] utilize residual vector quantization (RVQ) with reconstruction objectives. These methods achieve high-quality audio compression but may not capture higher-level semantic structures effectively.
Recognizing the complementary nature of semantic and acoustic tokens, recent works have attempted to combine these approaches. AudioLM [7] introduced a three-stage model: semantic modeling, coarse acoustic modeling, and fine acoustic modeling. While comprehensive, this approach introduces complexity and computational overhead. AudioPalm [32] further demonstrated the critical importance of large-scale training data and model parameters for effective multi-stage modeling, highlighting potential generalization issues in low-resource scenarios. An alternative hybrid approach, proposed by Zhang et al. [46], attempts to distill semantic information into acoustic tokens during RVQ model training. However, this method still requires additional pretraining and does not fully achieve a single-stage model architecture.
Despite these advancements, several challenges persist in the field of speech tokenization: i) balancing semantic and acoustic information in a unified representation; ii) reducing model complexity and computational requirements; iii) improving generalization to low-resource / out-of-domain data e.g. with mixed speech from multiple speakers, or multiple languages, or changing characteristics of recording equipment/sampling rate etc.; iv) developing truly single-stage tokenizers. Our proposed method, dMel, addresses these challenges by offering a training-free speech tokenization approach. By directly discretizing log mel-filterbanks into bins, it inherently preserves both semantic and acoustic information in a unified representation, while significantly reducing computational complexity.
Table 11: Results of ASR and TTS jointly trained model on LibriSpeech 960h.
ASR, WER↓ (%) TTS
test-clean test-other WER↓ (%) CER↓ (%)
@@ 4 Related Work
Modeling speech and text jointly is a challenging task, as speech signals are continuous and while text is discrete. Existing works have explored various approaches to address this challenge, including usage of separate encoders for different modalities [1, 5]. Bai et al. [3] proposed an encoder-only model A3T for speech-text modeling, by introducing alignment embedding to encourage cross-modal transfer between text and speech. Although A3T achieved good performance on speech synthesis and editing tasks, it cannot generate text and cannot generalize to longform generation because of its encoder-only architecture and mask-reconstruction training strategy. VioLA [39] also targets a unified speech-text model which can generate speech and text with a single model, but it is specifically designed for the Encodec [11] style feature, and compelled to model speech tokens in a multi-stage hierarchical manner. Maiti et al. [24] proposed a LM-style model VOXTLM, to model speech and text jointly. However, VOXTLM is only models the HuBERT semantic tokens, and relies on an external generation model to transform semantic tokens into waveform, but the speaker and acoustic information are lost. In comparison, the model architecture in this paper is a simple, single stage LM-style transformer model, and can handle both the speech generation and text generation tasks.